# Seo Hyun Kim1     Kai Tzu-iunn Ong1 Taeyoon Kwon1     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1

### v1

@@ 3.2 Baselines
To evaluate Theanine in long-term conversation, besides naive settings using all past dialogues or memories, we include the following baselines.
Memory retrieval. Following Xu et al. (2022a) and Xu et al. (2022b), this setting uses a retriever to retrieve memories with the current dialogue context to augment response generation.
+ Memory update. We use LLMs to perform the memory updating mechanism developed by Bae et al. (2022) to update the memory pool at the end of each session to enhance the above baselines.
RSum-LLM. Proposed by Wang et al. (2023), it uses an LLM to (i) recursively summarize and update the memory pool at the end of each session, (ii) select top memories based on the current dialogue, and (iii) generate responses.
@@
COMEDY. Proposed by Chen et al. (2024b), rather than memory retrieval, it first summarizes session-level past summaries and then compresses all of them into short sentences of events, user portraits (behavior patterns, emotion, etc.) and user-bot relation, using them to augment response generation.
@@ 3.3 Models and Implementation Details
Large language models. In all experiments including baselines, we adopt ChatGPT (gpt-3.5-turbo-0125) from OpenAI (2023) for (i) memory summarization, (ii) memory update, and (iii) response generation. We set the temperature to $0.75$.
@@
Retrievers. We use text-embedding-3-small (OpenAI, 2024b) from OpenAI to calculate the similarity of texts. For the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II Preparation), we set $j$ and $k$ to 3. For the “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al. (2022a).
Dialogue sessions. We perform evaluation using sessions 3-5 from both datasets, as all settings are almost identical before the end of session 2 (there is no memory to update before then).
4 Evaluation Scheme 1: Machine and Human Evaluation
To evaluate [Uncaptioned image]Theanine, we first start with G-Eval (Liu et al., 2023), an LLM-based framework for assessing LMs’ responses, and human evaluation. [^7] Details for G-Eval and human evaluation are in Appendix D and F. Dataset-specific results are in Appendix E. Here, we focus on the following research questions (RQs):
RQ1. Can Theanine generate context-rich responses, rather than those generic and plain?
RQ2. Can Theanine reference past conversations better than other methods?
RQ3. Do phases in Theanine make sense?
RQ4. Do humans agree with Theanine’s procedures, and how do they think of the final responses?
[Uncaptioned image]Theanine yields responses that are more detail-rich (RQ1).
In Table 1, we first assess whether generated responses are detail-rich, rather than generic and plain, i.e., specificity (Spe). In both datasets, Theanine yields significantly high win rates in the head-to-head comparison with all baselines except for RSum-LLM. We presume it is due to the fact that RSum-LLM tends to generate much longer responses (avg. 77 words per response) than other settings (ours = 32 words; avg of other baselines = 30 words).
[Uncaptioned image]Theanine utilizes past memories efficiently and correctly (RQ2).
In long-term conversations, it is crucial and necessary to further assess (i) whether the dialogue system utilizes information from the past when necessary (i.e., memorability) and make sure (ii) the response does not contradict previous conversations (i.e., consistency). We present the results in Table 1. Theanine demonstrates high win rates in the comparisons with baselines. However, COMEDY performs almost as well as us regarding memorability, which we assume is because it always applies two additional LLMs to extract and compress past conversations into concise memos before each response generation (no retrieval). However, as we expected, always referencing concise and recent user information can lead to contradiction to older information, thus clearly beaten by Theanine in terms of consistency (Cst).
Ablation results: Phases in [Uncaptioned image]Theanine matter (RQ3).
Table 2 presents the results where key phases in Theanine are removed. We observe that the removal of relation-aware memory linking (Phase I-2) and context-aware timeline refinement (Phase II-2) leads to large performance drops regarding the reference of past information (i.e., the original Theanine shows higher win rates), especially when timeline refinement is absent.
This justifies our methodological designs: (i) In graph construction, linking memories with cause-effect relation benefits RG more than naive connection using only text similarity and chronological order; (ii) By further refining memories built off-line (i.e., without considering current conversations) before RG, we can elicit contextual cues that are tailored for the current conversation.
Multi-session Chat (Session 3-5)
[Uncaptioned image]Theanine Ours’ Win Rate
          vs. Mem Cst Avg.
w/o Relation-aware Linking 62% 65% 64%
w/o Timeline Refinement 62% 69% 65%
Conversation Chronicles (Session 3-5)
w/o Relation-aware Linking 60% 65% 63%
w/o Timeline Refinement 64% 68% 66%
Table 2: Ablation results (p-values $<0.001$).
Figure 5: Results of human evaluation.
Humans think [Uncaptioned image]Theanine makes sense and properly references past conversations.
As reported in Figure 5, human judges largely agree (92%) that Theanine’s relation-aware memory linking is performed appropriately, and they all agree (100%) that timeline refinement can elicit more helpful/tailored information for response generation. Most importantly, Theanine produces responses that either entail (i.e., faithfully reflect) past conversations (34%) or, at worse, neutral statements (57%) in 91% of the cases from human standpoints, manifesting its’ efficacy in referencing past information in long-term conversations.
Examples of [Uncaptioned image]Theanine in terms of each phase are in Appendix G.
@@ 4 Evaluation Scheme 1: Machine and Human Evaluation
Despite the success of G-Eval (Liu et al., 2023) in evaluating responses of LMs, accessing dialogue systems’ ability to utilize past memories in long-term conversation still remains challenging. This is due to the unavailability of ground-truth mapping between current conversations and the correct memories for retrieval. Although we can resolve this by feeding the evaluator LLM (e.g., GPT-4) the entire memory pool or past history and prompt it to determine whether a response correctly recalls past conversations, the evaluation can be limited by the performance of the evaluator LLM itself.
To overcome this, along with Theanine, we present TeaFarm, a counterfactual-driven pipeline for evaluating memory-augmented response generation in long-term conversations.
5.1 Testing Dialogue Systems’ Memory via Counterfactual Questions
In TeaFarm, we proceed to “trick” the dialogue systems into generating incorrect responses, and the system must correctly reference past conversations to avoid being misled by us. Specifically, we talk to the dialogue system while acting as if a non-factual statement is true (thus counterfactual). Figure 6 presents some examples of counterfactual questions and the corresponding facts.
@@ 5.1 Testing Dialogue Systems’ Memory via Counterfactual Questions
In practice, when we want to evaluate a system that has been interacting with the user for multiple sessions, we first (1) collect all past conversations and summarize them session by session. Then, we (2) feed the LLM [^8] We apply GPT-4 (gpt-4) with a temperature of $0.75$. the collected summaries in chronological order such that it can capture the current stage of each discussed event, e.g., “Speaker B has never been to Japan at this moment”, and (3) generate counterfactual questions from the perspective of both speakers (and the correct answers). After that, we (4) kick off a new conversation, chat for a while, then (5) naturally ask the counterfactual question, and (6) assess the correctness of its response. Overview illustration and prompts for TeaFarm are in Appendix B and H.
@@ 5.2 TeaFarm Results
[Uncaptioned image]Theanine 0.17 0.24 0.21
      w/o Relation-aware Linking 0.06 0.13 0.10
      w/o Timeline Refinement 0.15 0.16 0.16
Table 3: TeaFarm results regarding correct/total percentages of model responses. Tested in conversations from MSC and CC with 200 counterfactual questions.
In Table 3, we present model performance in TeaFarm evaluations. Theanine outperforms all baselines, especially in CC. Ablations perform worse than original Theanine, again proving the efficacy and necessity of relation-aware linking and timeline refinement. Surprisingly, all settings result in low correct rates, qualifying TeaFarm as a proper challenge for stress-testing dialogue systems in long-term conversations.
Interestingly, baselines that utilize retrievers (just like Theanine) show superior performance than those relying on LLMs for memory selection (i.e., RSum-LLM, MemoChat, and COMEDY). [^9] Memory update does not affect Memory Retrieval’s performance. We believe it is because the counterfactual questions are made to counter the newest stage of each event. The removal of older memories thus does not have much impact. This pattern, unexpectedly, supports our idea of revisiting memory retrieval in the era of LLMs.
5.3 Dataset for TeaFarm: [Uncaptioned image]TeaBag
As a byproduct of TeaFarm, we curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of (1) 100 episodes of original conversations from MSC and CC (session 1-5), (2) follow-up conversations (session 6), where the speaker will naturally converse toward the counterfactual questions, and (3) the corresponding answers. TeaBag serve as a useful tool for stress-testing if a system can correctly reference past conversations. Dataset details are in Appendix C.
6 Further Discussions
@@
In Phase I-2, we choose to link $m_{new}$ to only the most recent associative memories in each $C_{i}$, rather than all associative memories in $M_{a}$. Figure 7 compares Theanine with Theanine-All, where $m_{new}$ is linked to all $m\in M_{a}$. We find that although Theanine-All costs 25% more, it is beaten by original Theanine. We believe this is because linking $m_{new}$ to all $M_{a}$ leads to more short timelines (e.g., an $m_{new}$ from session 4 is directly linked to an old memory from session 1), which provide less information than timelines built with relevant memories from more sessions. Regardless, it still yields lower loss rates than strong baselines. An open question here is whether we can reach a balance between cost and performance. We presume we can further improve the cost efficiency of Theanine in the long-term by training a memory linker using the collected memory-relation pairs from LLMs via knowledge distillation. We leave it to future work.
@@
Another inquiry left might be whether the growing span of memory would hinder memory retrieval in Theanine if there ever were hundreds of sessions. Although this can be a serious issue for conventional memory retrieval, we argue that it will be much mitigated in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines. This works as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements. Similarly, this also acts as a second insurance against mis-retrieved memories that are actually not helpful.
Figure 7: API cost and G-Eval results. Red shows baselines that lose even more than Theanine-All in the head-to-head comparisons with original Refer to captionTheanine.
@@ 7 Related Work
Since the release of MSC (Xu et al., 2022a), there have been many studies on long-term conversations. For instance, Bae et al. (2022) train a classifier as the memory updater to handle outdated memories in phone call scenarios. Recently, based on LLMs’ outstanding in-context learning ability, Wang et al. (2023) leverage LLMs to write, update, and select memories for response generation. Apart from LLMs’ power, human behaviors have also fostered methods for long-term conversations. For example, Zhong et al. (2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool for a longer time. Also drawing inspiration from how humans manage memories, our [Uncaptioned image]Theanine, on the other hand, addresses the connection between relevant events in both memory construction and retrieval, improving response quality in long-term conversations, which has never been explored to the best of our knowledge.
Memory-augmentation for personalized dialogue systems.
The trend of long-term interaction with dialogue systems promotes the adaptation of them for personalized needs (Chen et al., 2024a, c). As a pioneer, Xu et al. (2022b) train a persona extractor to create user-based memories in long-term conversations. However, training a personalized system for long-term use can be non-trivial due to the lack of data (Tseng et al., 2024). As a solution, Kim et al. (2024a) apply commonsense models and LLMs to augment existing long-term conversation datasets with high-quality persona sentences; Chen et al. (2024b) present a training-free LLM-based framework that extracts user behaviors from past conversations for personalized responses. Upon the success of LLMs, [Uncaptioned image]Theanine leverages them to build memory timelines, which represent the development of user personas/behaviors throughout the interaction, providing rich cues for personalized responses.
@@ 8 Conclusions
This work revisits memory management in multi-session/long-term conversations in the era of LLMs. Following how humans store and retrieve relevant memories as a whole, we present [Uncaptioned image]Theanine, a framework of timeline-augmented response generation, which elicits responses that are detail-rich and, more importantly, able to correctly reference previous conversations when necessary. We also propose a novel evaluation pipeline (TeaFarm) that addresses the limitation of G-Eval in long-term conversations as well as the dataset for it ([Uncaptioned image]TeaBag). We expect this paper to serve as a groundwork and inspiration for exploring more paradigms of memory construction/retrieval in dialogue systems.
@@ 9 Limitations
This work has the following limitations: First, the maximum amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets of multi-session conversations. Hence, Theanine’s performance in longer conversations has not been investigated. As we mentioned in Section 6, we presume that its’ effectiveness can still hold true to some degree even in such scenarios. Yet, we do acknowledge the need to apply additional modules that directly address the growing span of dialogue history/memories. One possible way is to introduce the summarize-then-compress paradigm in COMEDY (Chen et al., 2024b) to Theanine, which further compresses session-level summaries into a combined short user/event description.
Second, retrieving and refining timelines for response generation can be less computationally efficient than conventional memory retrieval, as there might exist scenarios where timelines are not necessary, e.g., when there is no need to reference past conversations. We believe this can be mitigated by incorporating an additional selection logic/module to choose between memory retrieval, timeline retrieval, and no retrieval at all.
Lastly, although we include many recent frameworks as baselines, we failed to compare Theanine with MemoryBank (Zhong et al., 2024), a framework inspired by Ebbinghaus’s forgetting curve. This is because the time intervals between sessions in MSC and CC are either mostly measured in hours or not clearly specified (e.g., “a few months later”), whereas MemoryBank requires precise time intervals in days to apply the forgetting curve. Also, data used for MemoryBank focuses on Chinese clinical scenarios, making it not feasible for our study. However, we remain positive about applying such a mechanism to improve Theanine in our ongoing research.

### v2

@@ 3.2 Baselines
To evaluate Theanine, besides naive baselines using all past dialogues or memories, we include:
Memory Retrieval. Following Xu et al. (2022a), we use a retriever to retrieve memories relevant to the current dialogue context to augment RG.
Memory Update. We use LLMs to run a widely used updating algorithm proposed by Bae et al. (2022) at the end of each session. The algorithm includes functionalities such as Change, Replace, Delete, Append, and more (see Appendix H).
RSum-LLM. An LLM-only generative method, which recursively summarizes and updates the memory pool and accordingly generates responses (w/o a retrieval module) (Wang et al., 2023).
@@
COMEDY. Proposed by Chen et al. (2024b), it uses LLMs to summarize session-level memories, compresses all of them into short events, user portraits (behavioral patterns, emotion, etc.) and user-bot relation, and selects them to augment RG.
@@ 3.3 Models and Implementation Details
Large language models. In all experiments including baselines, we adopt gpt-3.5-turbo-0125 (OpenAI, 2023) for (i) memory summarization (Table 4), (ii) memory update, and (iii) response generation. We set the temperature to $0.75$.
@@
Retrievers. We use text-embedding-3-small (OpenAI, 2024b) to calculate text similarity. In the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II), we set $j$ and $k$ to 3. For “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al. (2022a).
Dialogue sessions. We use sessions 3-5 of MSC and CC for evaluations, as all methods are almost the same in sessions 1-2 (no memory to update).
4 Evaluation Scheme 1: Automatic and Human Evaluations
To evaluate [Uncaptioned image]Theanine’s responses in long-term conversations, we follow common practices and conduct 3 types of evaluations: (i) Automatic evaluations; (ii) G-Eval (Liu et al., 2023), an LLM-based framework commonly used to evaluate LMs’ generation; (iii) human evaluation. We now present the main findings (details, prompts, and interfaces of evaluations in Scheme 1 are in Appendix E):
(Finding 1) Theanine outperforms baselines in response generation.
Table 1 reports agent performance in RG regarding both overlap-based and embedding-based metrics: Bleu-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Mauve (Pillutla et al., 2021), and BertScore (Zhang et al., 2020). Theanine, in both datasets, achieves superior response quality than diverse baselines. Despite slightly lower scores compared to Memory Retrieval in overlap-based metrics (i.e., B-4 and R-L) in MSC, Theanine largely outperforms it in terms of embedding-based metrics. Interestingly, including ours, methods w/o memory update generally yield higher scores, justifying our proposal towards an update-free memory management for lifelong dialogue agents.
Settings / Metrics B-4 R-L Mauve Bert
[Uncaptioned image]Theanine (Ours) 4.32 19.03 41.52 87.64
w/o Relation-aware Linking 4.07 18.58 39.69 87.57
w/o Timeline Refinement 4.03 18.82 41.34 87.66
Broken Down, Shuffled Timeline 4.15 18.70 38.49 87.61
Memory Retrieval 3.43 18.06 22.11 87.27
Table 2: Performance of our ablations (avg. of datasets).
(Finding 2 & 3) All phases contribute to performance; retrieving the timeline as a whole brings large improvement over conventional retrieval.
To gain deeper insights into our design, we show Theanine’s performance when relation-awareness in linking (Phase I-2) and Timeline Refinement (Phase II-2) are ablated. Also, to objectively assess whether Theanine’s retrieval (i.e., retrieving the timeline as a whole) improves retrieval quality, we include a setting where retrieved timelines are broken down into randomly ordered events such that retrieved memories during RG are in the same format as conventional top-$k$ retrieval.
In Table 2, we observe a ranking regarding contribution to performance: relation-aware linking $>$ retrieving timeline as a whole $>$ timeline refinement. We thus confirm the efficacy of constructing a memory graph with causal relations. Moreover, utilizing such a structure to collect timelines of relevant events yields better RG than conventional retrieval, despite the smaller $k$ (3 vs. 6) in initial retrieval. Refining timelines shows smaller performance gains, suggesting room for improvement in applying them for RG. We leave it to future work.
(Finding 4) Humans and G-Eval both reveal that Theanine retrieves memories that are more helpful for response generation.
Beyond agent responses, we further investigate how different memory construction methods affect retrieved memories’ helpfulness in augmenting RG. Given the same current dialogues as queries for retrieval, Figure 5 (left; G-Eval; right: Human Eval) shows head-to-head comparisons (ours vs baselines) regarding whose retrieved memories benefit response generation more. We observe higher win rates of Theanine in all comparisons, especially in human evaluations, suggesting that ours can facilitate a more helpful memory-augmentation for RG.
Figure 5: Human- (right) and machine-based (left) head-to-head comparisons between ours and baselines regarding the helpfulness of retrieved memories.
(Finding 5) Humans confirm that Theanine yields responses better entailing past interactions.
Now that the helpfulness of Theanine’s retrieved memories is validated, we proceed to investigate whether such helpful memories can help us facilitate reliable lifelong dialogue agents.
For that, we further ask a group of workers to specifically judge whether agent responses entail, contradict, or are neutral to the past via majority voting. In Figure 6, we find Theanine not only leading to a small number of contradictory responses (4%) but also showing the largest percentage (68%; out of 100) of responses that entail past conversations, significantly outperforming baselines. We argue that it is because our timeline-based approach elicits memories better at representing past interactions between speakers, thus leading to responses that more directly align with the past, which is important for dialogue agents to maintain long-term intimacy with users (Adiwardana et al., 2020). Also, such entail, non-contradictory nature of Theanine’s responses highlights its potential for applications in specialized domains, such as personalized agents for clinical scenarios, where entailment between agent responses and users’ past information (e.g., electrical health records or previous consulting sessions) is crucial for diagnostic decison-making (Tseng et al., 2024).
Figure 6: Human evaluations regarding to what extent the agent responses entail past conversations.
As a side note, Memory Update yields fewer contradictory responses (2%), indicating a potential trade-off between (i) removing outdated memories to prevent contradiction and (ii) preserving them to get richer information for RG (Kim et al., 2024a).
(Finding 6) Humans agree with Theanine’s intermediate procedures.
As reported in Figure 7, judges largely agree (92%) that Theanine properly assigns cause-effect relations to linked memories, explaining its contribution to performance. Also, they agree that timeline refinement successfully elicits more helpful information (100%; 100 samples in total) for RG. Examples of Theanine’s phases and RG are in Appendix G.
Figure 7: Human evaluation of our intermediate phases.
@@ 4 Evaluation Scheme 1: Automatic and Human Evaluations
Evaluating memory-augmented agents in long-term conversations is non-trivial due to the unavailability of ground-truth mapping between current conversations and correct memories for retrieval. Although we may resort to G-Eval by feeding evaluator LLMs (e.g., GPT-4) the entire past history and prompt it to determine whether a response correctly recalls the past, the evaluation can be largely limited by the performance of the evaluator LLM itself (Kim et al., 2024b).
To overcome this, along with Theanine, we present TeaFarm, a human-free counterfactual-driven pipeline for evaluating memory-augmented response generation in long-term conversations.
5.1 Testing Dialogue Agents’ Memory via Counterfactual Questions
In TeaFarm, we proceed to “trick” dialogue agents into generating incorrect responses, and agents must correctly reference past conversations to avoid being misled by us. Specifically, we talk to the dialogue agent while acting as if a non-factual statement is true (thus counterfactual). Figure 8 presents some examples of counterfactual questions and the corresponding facts.
@@ 5.1 Testing Dialogue Agents’ Memory via Counterfactual Questions
In practice (Figure 10), when we want to evaluate an agent that has been interacting with the user for sessions, we first (1) collect all past conversations and summarize them session by session. Then, we (2) feed a question generator LLM [^9] We apply GPT-4 (gpt-4) with a temperature of $0.75$. the collected summaries in chronological order such that it can capture the current stage of each discussed event, e.g., “Speaker B does not own a car”, and (3) generate counterfactual questions from the perspective of both speakers (and the correct answers). After that, we (4) kick off (i.e., simulate) a new dialogue session, chat for a while, then (5) naturally ask the counterfactual question, and (6) assess the correctness of its response. The overview figure, prompts, and synthesized data for TeaFarm are in Appendix CH, and D, respectively.
@@ 5.2 TeaFarm Results
[Uncaptioned image]Theanine 0.18 0.24 0.21
      w/o Relation-aware Linking 0.17 0.20 0.19
      w/o Timeline Refinement 0.16 0.19 0.18
Table 3: Success rates (SRs) of correctly recalling the past and not being fooled by the counterfactual questions in TeaFarm (tested with 200 questions).
In Table 3, Theanine shows higher SR than baselines, especially in CC. Ablations perform slightly worse than the original, again proving the efficacy of relation-aware linking and timeline refinement. Surprisingly, all settings have low SRs, qualifying TeaFarm as a proper pipeline for stress-testing dialogue agents in long-term conversations.
Interestingly, baselines using retrievers (same as Theanine) show superior performance than settings only relying on LLMs (i.e., RSum-LLM, MemoChat, and COMEDY). This, unexpectedly, supports our efforts in developing a new paradigm of memory management in the era of LLMs. [^10] Memory update does not affect Memory Retrieval’s performance. We believe it is because counterfactual questions are made to counter the newest stage of each event. The removal of older memories thus does not have much impact.
6 Further Analyses and Discussions
@@
A concern of Theanine is the API cost. Regardless, we argue that it is competitive when both performance and cost are taken into account. Figure 9 plots response quality (Mauve score) against the API cost. [^11] Calculated Based on session 5, which involves most memories for management. We use Mauve for its stronger correlation with humans over diverse metrics (Pillutla et al., 2021). We find Theanine and all ablations not only outperform all baselines but also lie on the Pareto frontier, indicating an efficient cost-performance trade-off. This suggests Theanine’s value when performance is prioritized over API costs.
Figure 9: Cost-performance comparisons.
@@
Another inquiry left might be whether the growing span of memory will eventually hinder memory retrieval in Theanine if there ever are hundreds of sessions. Although this can be a serious issue for conventional memory retrieval, we presume that it will be mitigated to some extent in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines. This serves as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements to further adjust retrieved timelines based on current dialogue context such that it provided tailored information for the response generator. This acts as a second insurance against sub-optimal retrieval.
@@ 7 Related Work
Since MSC, there are several studies on long-term conversations: Bae et al. (2022) train a classifier to update old memories in phone call scenarios. As we enter the era of LLMs, Wang et al. (2023) and Li et al. (2024) leverage LLMs to write and update memories for RG. Apart from LLMs’ power, human behaviors also foster methods in this field. For example, Zhong et al. (2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool longer. Recently, Maharana et al. (2024) also adopt the concept of timelines. However, in their work, a timeline is a fixed, pre-defined series of events (potentially unrelated), which simply serve as a user profile for synthesizing dialogue data. By contrast, in our work, a timeline is built with relevant events, which are dynamically linked and retrieved as the conversation goes on, serving as useful cues for response generation.
Memory-augmentation for personalized dialogue agents.
The trend of long-term interaction with autonomous agents promotes their adaptation for personalized needs (Chen et al., 2024a, c). As a pioneer, Xu et al. (2022b) train a persona extractor to create user-based memories. However, training personalized agents for long-term use can be non-trivial due to the lack of data (Tseng et al., 2024). As a solution, Kim et al. (2024a) apply commonsense models and LLMs to augment existing long-term data with high-quality persona sentences; Chen et al. (2024b) present a training-free LLM-based framework that extracts user behaviors from conversations for personalized RG. Upon the success of LLMs, Theanine leverages them to build memory timelines. These timelines represent the development of interactions and lead to responses that better entail speaker information, establishing Theanine’s potential for personalized agents.
@@ 8 Conclusions
This paper presents a first-ever timeline-based framework for memory-augmented dialogue agents in long-term conversations. Comprehensive evaluations show that [Uncaptioned image]Theanine facilitates more beneficial augmentation, leading to responses that are closer to ground-truths and more aligned with speakers’ past interactions. Its effectiveness is further confirmed in TeaFarm, a QA-based pipeline we propose to address the limitation of G-Eval and human efforts in assessing memory-augmented RG. We expect our novel approach to serve as a new foundation towards lifelong dialogue agents.
@@ 9 Limitations
First, the amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets. As we mentioned in Section 6, we presume that Theanine’s effectiveness can still hold true to some degree in longer conversations. Yet, we do acknowledge the need to apply additional modules that directly address the growing span of dialogue history/memories, such as introducing the summarize-then-compress paradigm in COMEDY (Chen et al., 2024b) to compress session-level summaries into a combined short user/event description.
Second, although we include many recent frameworks as baselines, we failed to compare Theanine with MemoryBank (Zhong et al., 2024), a framework inspired by Ebbinghaus’s forgetting curve. This is because the time intervals between sessions in MSC and CC are either mostly measured in hours or not clearly specified (e.g., “a few months later”), whereas MemoryBank requires precise time intervals in days to apply the forgetting curve. Also, data used for MemoryBank focuses on Chinese clinical scenarios, making it not feasible for our study. However, we remain positive about applying such a mechanism to improve Theanine in our ongoing research.
Lastly, API-based LLMs may introduce risks such as privacy issues. A possible solution is to apply Theanine to small open-source LMs for secure, local usage. While there exist challenges in data collection, one may achieve this by (i) collecting synthesized conversations with GPT-generated user profiles, (ii) running Theanine on these data, and (iii) using the outputs of each phase to train student LMs (i.e., distillation from teacher LLMs).